OpenStudy (anonymous):

Find x. cos(x-120) = sin2x Final answer in rad.

OpenStudy (tkhunny):

Have you considered expansion by identity? \(\cos(a-b) = \cos(a)\cos(b)+\sin(a)\sin(b)\) \(\sin(2x) = 2\sin(x)\cos(x)\)

OpenStudy (anonymous):

I got lost after \[- cosx - \sqrt{3}sinx = 4sinxcosx\]

OpenStudy (tkhunny):

We'll get back to the algebra in a minute. Let's check the trigonometry, first. \(\cos(x-120º) = \cos(x)\cos(2\pi/3) + \sin(x)\sin(2pi/3) = (-1/2)\cos(x) + (\sqrt{3}/2)\sin(x)\) Whoops! Looks like we missed a sign, in there. This gets us to where you were: \(-\cos(x) + \sqrt{3}\sin(x) = 4\sin(x)\cos(x)\)

OpenStudy (tkhunny):

Not really seeing how to go about that analytically. Everything has period \(2\pi\). That simplifies things. Various numerical methods can be employed to drag out the solutions. I get, for k an integer: \(x = 1.2217304764 + 2k\pi\) \(x = 3.3161255788 + 2k\pi\) \(x = 5.4105206812 + 2k\pi\) \(x = 5.7595865316 + 2k\pi\) Somewhat to my surprise, these could be: \(x = 7\pi/18 + 2k\pi\) \(x = 19\pi/18 + 2k\pi\) \(x = 31\pi/18 + 2k\pi\) \(x = 11\pi/6 + 2k\pi\) So then, how can we solve analytically? Ideas?